# Regular functional covering numbers

### v1

@@ Abstract
$\max\Bigl\{N(f,t\odot g),\,N(f^{\ast},t\odot g),\,N(g,t\odot f),\,N(g,t\odot f^{\ast})\Bigr\}\leqslant\exp\!\left(\frac{\gamma_{n}^{2}\,n}{t}\right),$
where $f^{\ast}$ denotes the Legendre dual of $f$, $(t\odot f)(x)=f(x/t)$ is the $t$-homothety of $f$, and $\gamma_{n}\leqslant c(\ln n)^{2}$. Our result shows that the isotropic position of a log-concave function already provides an almost $1$-regular functional $M$-position.
@@ 1 Introduction
for an absolute constant $\beta>0$, where $B_{2}^{n}$ is the Euclidean unit ball and $\widetilde{K}^{\circ}$ is the polar body of $\widetilde{K}$. Background material on convex bodies and log-concave functions is collected in Section 2.
@@
$\max\Bigl\{N(f,\,t\odot g),\,N(g,\,t\odot f^{\ast})\Bigr\}\;\leqslant\;\exp\!\left(\frac{\gamma_{n}^{2}\,n}{t}\right),$
@@
$\max\Bigl\{N(f^{\ast},\,t\odot g),\,N(g,\,t\odot f),\,\Bigr\}\;\leqslant\;\exp\!\left(\frac{\delta_{n}^{2}\,n}{t}\right),$
@@
We outline the main ideas of the proof. To each isotropic $f$ we associate the convex body
@@
One may compare $R_{f}$ with the isotropic convex body $K_{n+1}(f)$ introduced by K. Ball [7], and verify that their geometric distance is bounded by an absolute constant. We then use an observation of the second named author [28], based on E. Milman’s sharp $M^{\ast}$-estimate [22] and the recent optimal $M$-estimate of Bizeul and Klartag [9] for isotropic convex bodies, to show that $R_{f}$ satisfies almost $2$-regular covering estimates. These geometric bounds are transferred to the functional covering numbers $N(f,\,t\odot g)$ and $N(g,\,t\odot f)$ using a decomposition of $\exp(-50n\|\cdot\|_{R_{f}})$, a corresponding decomposition of the Gaussian $g$, and basic structural properties of functional covering numbers.
@@
Let $f:\mathbb{R}^{n}\to[0,\infty)$ be an isotropic geometric log-concave function.. Then,
@@
The definition of the $\mathcal{A}$-transform appears in the book by Rockafellar [27, page 136], where it is also proved that it commutes with the Legendre transform. However, the polarity transform was introduced and studied in depth by Artstein-Avidan and Milman in [1] as the functional extension of convex-body polarity and plays a central role in functional versions of the Blaschke-Santaló and Bourgain-Milman inequalities.
@@
$N(g,\,t\odot f_{\mathcal{A}})\;\leqslant\;\exp\!\left(\frac{\gamma_{n}^{2}\,n}{t}\right)\quad\text{and}\quad N(f_{\mathcal{A}},\,t\odot g)\;\leqslant\;\exp\!\left(\frac{\delta_{n}^{2}\,n}{t}\right),$
@@ 2 Convex bodies and log-concave functions
2.2. Log-concave functions. A function $f:\mathbb{R}^{n}\to[0,\infty)$ is called log-concave if it can be written in the form $f=e^{-\varphi}$, where $\varphi:\mathbb{R}^{n}\to(-\infty,\infty]$ is a proper, lower semi-continuous (l.s.c.) convex function. Properness means that $\mathrm{dom}(\varphi):=\{x\in\mathbb{R}^{n}:\varphi(x)<\infty\}\neq\varnothing$, and l.s.c. convexity ensures that $f$ is upper semi-continuous and satisfies the classical log-concavity inequality $f\big((1-\lambda)x+\lambda y\big)\geq f(x)^{1-\lambda}f(y)^{\lambda}$ for all $x,y\in\mathbb{R}^{n},\;\lambda\in(0,1)$.
@@
In view of the identity $\mathcal{L}(\varphi\square\psi)=\mathcal{L}\varphi+\mathcal{L}\psi$, the operation $f\star g$ allows us to define the analogue of Minkowski addition of log-concave functions.
@@
where
$\operatorname{Cov}(f):=\frac{1}{\int_{\mathbb{R}^{n}}f(x)\,dx}\int_{\mathbb{R}^{n}}x\otimes x\,f(x)\,dx-\operatorname{bar}(f)\otimes\operatorname{bar}(f)$
is the covariance matrix of $f$. A log-concave function $f$ is called isotropic if
$\operatorname{bar}(f)=0\quad\int_{\mathbb{R}^{n}}f(x)\,dx=1\quad\text{and}\quad\operatorname{Cov}(f)=\lambda_{f}^{2}I_{n}$
for some $\lambda_{f}>0$.
A convex body $K$ in $\mathbb{R}^{n}$ is called isotropic if $\mathrm{vol}_{n}(K)=1$, ${\rm bar}(K)=0$, and ${\rm Cov}(\mu_{K})=L_{\mu_{K}}^{2}I_{n}$, where $\mu_{K}$ is the uniform measure on $K$. Note that $K$ is isotropic if and only if its indicator function $\mathds{1}_{K}$ is isotropic.
It is straightforward to check that any centered log-concave function $f$ admits an invertible linear map $T\in GL_{n}$ such that $f_{1}:=f\circ T$ is isotropic; moreover, $L_{f_{1}}=L_{f}$. Since $f_{1}(0)=\|f_{1}\|_{\infty}$, dividing $f_{1}$ by $f_{1}(0)$ yields a function in $LC_{g}(\mathbb{R}^{n})$. From (2.1) we obtain
$\int_{\mathbb{R}^{n}}f_{1}(x)\,dx=a^{n},\qquad a:=\lambda_{f_{1}}/L_{f}.$
Define $f_{2}(x):=f_{1}(ax)$. Then
$f_{2}(0)=\|f_{2}\|_{\infty}=1,\qquad\int_{\mathbb{R}^{n}}f_{2}(x)\,dx=1,\qquad\operatorname{Cov}(f_{2})=L_{f}^{2}I_{n}.$
A log-concave function satisfying these properties is called an isotropic geometric log-concave function. Thus every centered log-concave $f$ with finite positive integral admits an isotropic position $\tilde{f}=f\circ T$ for some $T\in GL_{n}$.
Bourgain’s slicing problem [10] asked if there exists an absolute constant $C>0$ such that
(2.2) $L_{n}:=\max\{L_{K}:K\ \hbox{is an isotropic convex body in}\ \mathbb{R}^{n}\}\leqslant C.$
@@
where the supremum is taken over isotropic log-concave functions $f$ on $\mathbb{R}^{n}$. Bourgain [11] showed that $L_{n}\leqslant c\,n^{1/4}\ln n$, improved by Klartag [20] to $L_{n}\leqslant c\,n^{1/4}$. These were the best known bounds until 2020. In a breakthrough, Chen [13] proved that for every $\epsilon>0$,
@@ E. Milman [22] proved that if $K$ is isotropic then
The dual estimate was recently obtained by Bizeul and Klartag [9]:
@@
Let $r_{K}$ denote the radius of the Euclidean ball with the same volume as $K$, i.e. $\mathrm{vol}_{n}(K)=\mathrm{vol}_{n}(r_{K}B_{2}^{n})$. For isotropic bodies we have $\mathrm{vol}_{n}(K)=1$ and hence this radius depends only on the dimension: $r_{K}=r_{n}=\omega_{n}^{-1/n}$, and $r_{n}\simeq\sqrt{n}$.
@@
These estimates provide a quantitative control of covering numbers for isotropic bodies and their polars, which is crucial in applications. In particular, they will play an essential role in the proof of Theorem 1.3.
@@
Let $f,g\in LC_{g}((\mathbb{R}^{n})$. Then, $M(f,\overline{g})=N(f,g)$, where $\overline{g}(x)=g(-x)$.
@@ 3 Regular functional covering numbers
Let $f:\mathbb{R}^{n}\to[0,\infty)$ be a centered log-concave function with $f(0)>0$. We associate to $f$ two classical families of convex bodies, denoted by $\{R_{t}(f)\}_{t\geqslant 1}$ and $\{K_{t}(f)\}_{t\geqslant 1}$. First, for every $t\geqslant 1$, define
@@
The second family of convex bodies $K_{t}(f)$ was introduced by K. Ball, who also proved their convexity in [7]. For every $t\geqslant 1$, define
@@
Its radial function is given by
@@
For $0<t\leqslant s$ one has the inclusions (see [12, Prop. 2.5.7])
@@
Let $f:\mathbb{R}^{n}\to[0,\infty)$ be a centered log-concave function with $f(0)>0$. For every $s\geqslant t$ we have
@@
(3.7) $r_{f}=(\mathrm{vol}_{n}(R_{f}))/\omega_{n})^{1/n}.$
@@
(3.8) $\max\Bigl\{N(R_{f},\,tr_{f}B_{2}^{n}),\,N(B_{2}^{n},\,tr_{f}(R_{f})^{\circ})\Bigr\}\leqslant\exp\left(\tfrac{c_{1}\gamma_{n}^{2}n}{t^{2}}\right),$
for some $\gamma_{n}\geqslant 1$ and every $t\geqslant 1$, then
$N(f,t\odot g)\leqslant\exp\!\left(\frac{\gamma_{n}^{2}n}{t}\right)\qquad\text{for all }t\geqslant 1.$
@@
If $x\in R_{f}$, then $f(x)\leqslant 1=\mathds{1}_{R_{f}}(x)$. If $x\notin R_{f}$, then $\|x\|_{R_{f}}>1$ and by definition $f(x/\|x\|_{R_{f}})=\exp(-50n)$. Log-concavity yields
$f(x)\leqslant\exp\!\left(-50n\,\|x\|_{R_{f}}\right).$
@@
(3.10) $\exp(-50n\|x\|_{R_{f}})=\sum_{k=0}^{\infty}e^{-50n\|x\|_{R_{f}}}\,\mathds{1}_{\{k\leq 50n\|x\|_{R_{f}}<k+1\}}(x)\leqslant\sum_{k=0}^{\infty}e^{-k}\,\mathds{1}_{\frac{k+1}{50n}R_{f}}(x).$
@@
Since $\mathrm{vol}(R_{f})^{1/n}\approx 1$, we have $r_{f}\approx\sqrt{n}$, and for $x\in\sqrt{t}r_{f}B_{2}^{n}$,
$\exp\!\left(-\frac{|x|^{2}}{2t^{2}}\right)\geqslant\exp\left(-\frac{r_{f}^{2}}{2t}\right)\geqslant\exp(-cn/t),$
@@
$\mathds{1}_{\sqrt{t}r_{f}B_{2}^{n}}(x)\leqslant e^{cn/t}(t\odot g)(x),$
@@
(3.14) $N(\mathds{1}_{\sqrt{t}r_{f}B_{2}^{n}},t\odot g)\leqslant e^{cn/t}.$
@@
(3.15) $N(\mathds{1}_{R_{f}},t\odot g)\leqslant\exp\!\left(\frac{c_{1}\gamma_{n}^{2}n}{t}\right).$
@@
$N(\mathds{1}_{\frac{k+1}{50n}R_{f}},t\odot g)$ $\leqslant N(\mathds{1}_{\frac{k+1}{50n}R_{f}},\mathds{1}_{\sqrt{t}r_{f}B_{2}^{n}})\,N(\mathds{1}_{\sqrt{t}r_{f}B_{2}^{n}},t\odot g)$
$\leqslant e^{cn/t}N(R_{f},\sqrt{t}r_{f}B_{2}^{n})\,N\!\left(B_{2}^{n},\tfrac{50n}{k+1}B_{2}^{n}\right)\,.$
@@
$N(\mathds{1}_{\frac{k+1}{50n}R_{f}},t\odot g)\leqslant\exp\!\left(\frac{c_{2}\gamma_{n}^{2}n}{t}\right)e^{\frac{k+1}{25}}.$
@@
$\sum_{k=0}^{\infty}e^{-k}N(\mathds{1}_{\frac{k+1}{50n}R_{f}},t\odot g)\leqslant C_{1}\exp\!\left(\frac{c_{2}\gamma_{n}^{2}n}{t}\right),$
@@
$N(f,t\odot g)\leqslant\exp\!\left(\frac{c_{1}\gamma_{n}^{2}n}{t}\right)+C_{1}\exp\!\left(\frac{c_{2}\gamma_{n}^{2}n}{t}\right),$
@@
(3.16) $\max\Bigl\{N(r_{f}B_{2}^{n},\,tR_{f}),\,N(r_{f}R_{f}^{\circ},\,tB_{2}^{n})\Bigr\}\leqslant\exp\!\left(\tfrac{c_{1}\delta_{n}^{2}n}{t^{2}}\right)$
for some $\delta_{n}\geqslant 1$ and every $t\geqslant 1$, then
$N(g,t\odot f)\leqslant\exp\!\left(\frac{\delta_{n}^{2}n}{t}\right)\qquad\text{for all }t\geqslant 1.$
@@
Let $t\geqslant 1$. If $x\in R_{f}$, then by definition $f(x)\geqslant\exp(-50n)$, and log-concavity gives
$f(x/t)\geqslant f(x)^{1/t}\geqslant\exp(-50n/t).$
Hence
@@
and Lemma 2.2 (i) yields
(3.17) $N(g,t\odot f)\leqslant e^{-50n/t}\,N(g,\mathds{1}_{R_{f}}).$
@@
$g(x)=\sum_{k=0}^{\infty}g(x)\,\mathds{1}_{\{akr_{f}\leqslant|x|<a(k+1)r_{f}\}}(x)\leqslant\sum_{k=0}^{\infty}e^{-a^{2}k^{2}r_{f}^{2}/2}\,\mathds{1}_{a(k+1)r_{f}B_{2}^{n}}(x),$
@@
$\leqslant N\!\left(a(k+1)r_{f}B_{2}^{n},\,\tfrac{1}{\sqrt{t}}r_{f}B_{2}^{n}\right)\,N(r_{f}B_{2}^{n},\,\sqrt{t}R_{f}).$
@@
$N\!\left(a(k+1)r_{f}B_{2}^{n},\,\tfrac{1}{\sqrt{t}}r_{f}B_{2}^{n}\right)\leqslant\bigl(2\sqrt{t}\,a(k+1)+1\bigr)^{n},$
@@
$N(r_{f}B_{2}^{n},\sqrt{t}R_{f})\leqslant\exp\!\left(\frac{c_{1}\delta_{n}^{2}n}{t}\right).$

### v2

# arXiv:2512.04301v2[math.MG] 02 Mar 2026

@@ Abstract
$\max\Bigl\{N(f,t\odot g),N(g,t\odot f^{\ast})\Bigr\}\leqslant\exp\!\left(\gamma_{n}n/t\right)\quad\text{and}\quad\max\Bigl\{N(f^{\ast},t\odot g),N(g,t\odot f)\Bigr\}\leqslant\exp\!\left(\delta_{n}n/t\right),$
where $f^{\ast}$ denotes the Legendre dual of $f$, $(t\odot f)(x)=f(x/t)$ is the $t$-homothety of $f$, and $\gamma_{n}\leqslant c(\ln n)^{2}$, $\delta_{n}\leqslant c\ln n$. Our result shows that the isotropic position of a log-concave function already provides an almost $1$-regular functional $M$-position.
@@ 1 Introduction
for an absolute constant $\beta>0$, where $B_{2}^{n}$ is the Euclidean unit ball and $\widetilde{K}^{\circ}$ is the polar body of $\widetilde{K}$. Background material on convex bodies is collected in Section 2.
@@
$\max\Bigl\{N(f,\,t\odot g),\,N(g,\,t\odot f^{\ast})\Bigr\}\;\leqslant\;\exp\!\left(\frac{\gamma_{n}n}{t}\right),$
@@
$\max\Bigl\{N(f^{\ast},\,t\odot g),\,N(g,\,t\odot f),\,\Bigr\}\;\leqslant\;\exp\!\left(\frac{\delta_{n}n}{t}\right),$
@@
We outline the main ideas of the proof. To each isotropic geometric log-concave function $f$ we associate the convex body
@@
One may compare $R_{f}$ with the almost isotropic convex body $K_{n+1}(f)$ introduced by K. Ball [7], and verify that their geometric distance is bounded by an absolute constant. We then use an observation of the second named author [28], based on E. Milman’s sharp $M^{\ast}$-estimate [22] and the recent optimal $M$-estimate of Bizeul and Klartag [9] for isotropic convex bodies, to show that $R_{f}$ satisfies almost $2$-regular covering estimates. These geometric bounds are transferred to the functional covering numbers $N(f,\,t\odot g)$ and $N(g,\,t\odot f)$ using a decomposition of $\exp(-50n\|\cdot\|_{R_{f}})$, a corresponding decomposition of the Gaussian $g$, and basic structural properties of functional covering numbers.
@@
The definition of the $\mathcal{A}$-transform appears in the book by Rockafellar [27, page 136], where it is also proved that it commutes with the Legendre transform. However, the polarity transform was introduced and studied in depth by Artstein-Avidan and Milman in [1] as the functional extension of convex-body polarity and plays a central role in functional versions of the Blaschke–Santaló and Bourgain–Milman inequalities.
@@
$N(g,\,t\odot f_{\mathcal{A}})\;\leqslant\;\exp\!\left(\frac{\gamma_{n}n}{t}\right)\quad\text{and}\quad N(f_{\mathcal{A}},\,t\odot g)\;\leqslant\;\exp\!\left(\frac{\delta_{n}n}{t}\right),$
@@ 2 Convex bodies and log-concave functions
2.2. Log-concave functions. A function $f:\mathbb{R}^{n}\to[0,\infty)$ is called log-concave if it can be written in the form $f=e^{-\varphi}$, where $\varphi:\mathbb{R}^{n}\to(-\infty,\infty]$ is a proper, lower semi-continuous (l.s.c.) convex function. Properness means that $\mathrm{dom}(\varphi):=\{x\in\mathbb{R}^{n}:\varphi(x)<\infty\}\neq\varnothing$, and l.s.c. convexity ensures that $f$ is upper semi-continuous and satisfies the classical log-concavity inequality $f\big((1-\lambda)x+\lambda y\big)\geqslant f(x)^{1-\lambda}f(y)^{\lambda}$ for all $x,y\in\mathbb{R}^{n},\;\lambda\in(0,1)$.
@@
In view of the identity $\mathcal{L}(\varphi\square\psi)=\mathcal{L}\varphi+\mathcal{L}\psi$, the operation $f\star g$ allows us to define the analogue of Minkowski addition for log-concave functions.
@@
where $\operatorname{Cov}(f)$ is the covariance matrix of $f$ with entries
$\textrm{Cov}(f)_{i,j}:=\frac{\int_{\mathbb{R}^{n}}x_{i}x_{j}f(x)\,dx}{\int_{\mathbb{R}^{n}}f(x)\,dx}-\frac{\int_{\mathbb{R}^{n}}x_{i}f(x)\,dx}{\int_{\mathbb{R}^{n}}f(x)\,dx}\frac{\int_{\mathbb{R}^{n}}x_{j}f(x)\,dx}{\int_{\mathbb{R}^{n}}f(x)\,dx}.$
In this article, a log-concave function $f$ is called isotropic if
$\operatorname{bar}(f)=0,\quad\int_{\mathbb{R}^{n}}f(x)\,dx=1\quad\text{and}\quad\operatorname{Cov}(f)=\lambda_{f}^{2}I_{n}$
for some $\lambda_{f}>0$. We say that $f$ is an isotropic geometric log-concave function if it also satisfies $f(0)=\|f\|_{\infty}=1$.
A convex body $K$ in $\mathbb{R}^{n}$ is called isotropic if $\mathrm{vol}_{n}(K)=1$, ${\rm bar}(K)=0$, and ${\rm Cov}(\mu_{K})=L_{\mu_{K}}^{2}I_{n}$, where $\mu_{K}$ is the uniform measure on $K$. Note that $K$ is isotropic if and only if its indicator function $\mathds{1}_{K}$ is an isotropic geometric log-concave function.
It is straightforward to check that any centered log-concave function $f$ admits an invertible linear map $T\in GL_{n}$ such that the covariance matrix of $f_{1}:=f\circ T$ is a multiple $\lambda_{f_{1}}^{2}I_{n}$ of the identity; moreover, $L_{f_{1}}=L_{f}$. Dividing $f_{1}$ by $\|f_{1}\|_{\infty}$ yields a centered log-concave function $f_{2}$ with
$\|f_{2}\|_{\infty}=1,\quad\int_{\mathbb{R}^{n}}f_{2}(x)\,dx=a^{n},\quad a:=\lambda_{f_{1}}/L_{f}.$
Define $f_{3}(x):=f_{2}(ax)$. Then
$\|f_{3}\|_{\infty}=1,\qquad\int_{\mathbb{R}^{n}}f_{3}(x)\,dx=1,\qquad\operatorname{Cov}(f_{3})=L_{f}^{2}I_{n}.$
A log-concave function satisfying these properties is called a normalized isotropic log-concave function. Thus every centered log-concave $f$ with finite positive integral admits a normalized isotropic position $\tilde{f}=f\circ T$ for some $T\in GL_{n}$. If we make the additional assumption that $f(0)=\|f\|_{\infty}$ then we have that $\tilde{f}\in LC_{g}(\mathbb{R}^{n})$, i.e. $\tilde{f}$ is an isotropic geometric log-concave function.
Bourgain’s slicing problem [10] asks if there exists an absolute constant $C>0$ such that
(2.2) $L_{n}:=\max\{L_{\mu_{K}}:K\ \hbox{is an isotropic convex body in}\ \mathbb{R}^{n}\}\leqslant C.$
@@
where the supremum is taken over all isotropic log-concave functions $f$ on $\mathbb{R}^{n}$. Bourgain [11] showed that $L_{n}\leqslant c\,n^{1/4}\ln n$, improved by Klartag [20] to $L_{n}\leqslant c\,n^{1/4}$. These were the best known bounds until 2020. In a breakthrough, Chen [13] proved that for every $\epsilon>0$,
@@ E. Milman [22] proved that if $K$ is isotropic then
The dual estimate was recently obtained by Bizeul and Klartag [9]: if $K$ is isotropic, then
@@
Let $r_{K}$ denote the radius of the Euclidean ball with the same volume as $K$, i.e. $\mathrm{vol}_{n}(K)=\mathrm{vol}_{n}(r_{K}B_{2}^{n})$. For isotropic bodies we have $\mathrm{vol}_{n}(K)=1$ and hence this radius depends only on the dimension: $r_{K}=r_{n}=\omega_{n}^{-1/n}$, and $r_{n}\approx\sqrt{n}$.
@@
These estimates provide a quantitative control of covering numbers for isotropic convex bodies and their polars, which is crucial in applications. In particular, they will play an essential role in the proof of Theorem 1.3.
@@
Let $f,g\in LC_{g}(\mathbb{R}^{n})$. Then, $M(f,\overline{g})=N(f,g)$, where $\overline{g}(x)=g(-x)$.
@@ 3 Regular functional covering numbers
Let $f:\mathbb{R}^{n}\to[0,\infty)$ be a centered log-concave function with $f(0)>0$. We associate to $f$ two classical families of convex bodies, denoted by $\{R_{t}(f)\}_{t>0}$ and $\{K_{t}(f)\}_{t>0}$. First, for every $t>0$, define
@@
The second family of convex bodies $K_{t}(f)$ was introduced by K. Ball, who also proved their convexity in [7]. For every $t>0$, define
@@
The radial function of $K_{t}(f)$ is given by
@@
For $0<t\leqslant s$ one has the inclusions (see [12, Proposition 2.5.7])
@@
Let $f:\mathbb{R}^{n}\to[0,\infty)$ be a centered log-concave function with $f(0)>0$. For every $s\geqslant t>0$ we have
@@
Remark 3.4.
If $f$ is isotropic, then the convex body $K_{n+1}(f)$ is almost isotropic in the sense of [12, Definition 2.5.11]: for any $T\in GL_{n}$ such that $T(K_{n+1}(f))$ is isotropic, we have that
$\alpha^{-1}B_{2}^{n}\subseteq T(B_{2}^{n})\subseteq\alpha B_{2}^{n},$
where $\alpha\geqslant 1$ is an absolute constant. For a proof of this claim, see [12, Proposition 2.5.12]. It follows that
$N(K_{n+1}(f),tB_{2}^{n})=N(T(K_{n+1}(f)),tT(B_{2}^{n}))\leqslant N(T(K_{n+1}(f)),\alpha^{-1}tB_{2}^{n})$
for every $t\geqslant 1$, and similar inequalities are valid for the other three covering numbers $N(B_{2}^{n},tK_{n+1}(f))$, $N((K_{n+1}(f))^{\circ},tB_{2}^{n})$ and $N(B_{2}^{n},t(K_{n+1}(f))^{\circ})$. Therefore, (2.6) and (2.7) hold true for $K_{n+1}(f)$.
@@
(3.8) $r_{f}=\left(\frac{\mathrm{vol}_{n}(R_{f})}{\omega_{n}}\right)^{1/n}.$
@@
(3.9) $\max\Bigl\{N(R_{f},\,tr_{f}B_{2}^{n}),\,N(B_{2}^{n},\,tr_{f}(R_{f})^{\circ})\Bigr\}\leqslant\exp\left(\tfrac{c_{1}\kappa_{n}^{2}n}{t^{2}}\right),$
for some $\kappa_{n}\geqslant 1$ and every $t\geqslant 1$, then
$N(f,t\odot g)\leqslant\exp\!\left(\frac{\kappa_{n}n}{t}\right)\qquad\text{for all }t\geqslant 1.$
@@
If $x\in R_{f}$, then $f(x)\leqslant 1=\mathds{1}_{R_{f}}(x)$. If $x\notin R_{f}$, then $\lambda:=\|x\|_{R_{f}}>1$ and by definition $f(x/\lambda)=\exp(-50n)f(0)$. Log-concavity yields
$f(x/\lambda)\geqslant f(x)^{\frac{1}{\lambda}}f(0)^{1-\frac{1}{\lambda}}$
and hence
(3.11) $f(x)\leqslant\left(\frac{f(x/\lambda)}{f(0)}\right)^{\lambda}f(0)=\exp\!\left(-50n\,\|x\|_{R_{f}}\right)f(0)\leqslant\exp\!\left(-50n\,\|x\|_{R_{f}}\right).$
@@
(3.12) $\exp(-50n\|x\|_{R_{f}})=\sum_{k=0}^{\infty}e^{-50n\|x\|_{R_{f}}}\,\mathds{1}_{\{k\leqslant 50n\|x\|_{R_{f}}<k+1\}}(x)\leqslant\sum_{k=0}^{\infty}e^{-k}\,\mathds{1}_{\frac{k+1}{50n}R_{f}}(x).$
@@
Since $\mathrm{vol}(R_{f})^{1/n}\approx 1$, we have $r_{f}\approx\sqrt{n}$, and for $x\in\sqrt{\kappa_{n}t}r_{f}B_{2}^{n}$,
$\exp\!\left(-\frac{|x|^{2}}{2t^{2}}\right)\geqslant\exp\left(-\frac{\kappa_{n}r_{f}^{2}}{2t}\right)\geqslant\exp(-c\kappa_{n}n/t),$
@@
$\mathds{1}_{\sqrt{\kappa_{n}t}r_{f}B_{2}^{n}}(x)\leqslant e^{c\kappa_{n}n/t}(t\odot g)(x),$
@@
(3.16) $N(\mathds{1}_{\sqrt{\kappa_{n}t}r_{f}B_{2}^{n}},t\odot g)\leqslant e^{c\kappa_{n}n/t}.$
@@
(3.17) $N(\mathds{1}_{R_{f}},t\odot g)\leqslant\exp\!\left(\frac{c_{1}\kappa_{n}n}{t}\right).$
@@
$N(\mathds{1}_{\frac{k+1}{50n}R_{f}},t\odot g)$ $\leqslant N(\mathds{1}_{\frac{k+1}{50n}R_{f}},\mathds{1}_{\sqrt{\kappa_{n}t}r_{f}B_{2}^{n}})\,N(\mathds{1}_{\sqrt{\kappa_{n}t}r_{f}B_{2}^{n}},t\odot g)$
$\leqslant e^{c\kappa_{n}n/t}N(R_{f},\sqrt{\kappa_{n}t}r_{f}B_{2}^{n})\,N\!\left(B_{2}^{n},\tfrac{50n}{k+1}B_{2}^{n}\right)\,.$
@@
$N(\mathds{1}_{\frac{k+1}{50n}R_{f}},t\odot g)\leqslant\exp\!\left(\frac{c_{2}\kappa_{n}n}{t}\right)e^{\frac{k+1}{25}}.$
@@
$\sum_{k=0}^{\infty}e^{-k}N(\mathds{1}_{\frac{k+1}{50n}R_{f}},t\odot g)\leqslant C_{1}\exp\!\left(\frac{c_{2}\kappa_{n}n}{t}\right),$
@@
$N(f,t\odot g)\leqslant\exp\!\left(\frac{c_{1}\kappa_{n}n}{t}\right)+C_{1}\exp\!\left(\frac{c_{2}\kappa_{n}n}{t}\right),$
@@
Remark 3.6.
(i) In the proof of Proposition 3.5 we used only the bound $N(R_{f},\,tr_{f}B_{2}^{n})\leqslant\exp\left(\tfrac{c_{1}\kappa_{n}^{2}n}{t^{2}}\right)$ from (3.9). Moreover, in (3.11) we only use the fact that $f(0)\leqslant\|f\|_{\infty}=1$. Therefore, Proposition 3.5 holds for every log-concave function $f:\mathbb{R}^{n}\to[0,\infty)$ with $\|f\|_{\infty}=1$ that satisfies the first bound in (3.9).
(ii) Note also that, instead of decomposing $\exp(-50n\|x\|_{R_{f}})$ into a sum, as in (3.12), we may write
$e^{-50n\|x\|_{R_{f}}}=50n\int_{0}^{\infty}e^{-50ns}\mathds{1}_{sR_{f}}(x,s)\,ds=\int_{0}^{\infty}e^{-u}\mathds{1}_{\frac{u}{50n}R_{f}}(x,u)\,du$
and then estimate the functional covering number $N(\exp(-50n\,\|x\|_{R_{f}}),t\odot g)$ writing
$N(\exp(-50n\,\|x\|_{R_{f}}),t\odot g)\leqslant\int_{0}^{\infty}e^{-u}N(\mathds{1}_{\frac{u}{50n}R_{f}},t\odot g)\,du.$
Using the same argument as in the proof above, we get
$N(\mathds{1}_{\frac{u}{50n}R_{f}},t\odot g)\leqslant\exp\!\left(\frac{c_{2}\kappa_{n}n}{t}\right)e^{\frac{u}{25}},$
and hence
$N(\exp(-50n\,\|x\|_{R_{f}}),t\odot g)\leqslant\exp\!\left(\frac{c_{2}\kappa_{n}n}{t}\right)\int_{0}^{\infty}e^{-u}e^{\frac{u}{25}}du.$
The last integral is bounded by an absolute constant, and Proposition 3.5 follows.
@@
(3.18) $\max\Bigl\{N(r_{f}B_{2}^{n},\,tR_{f}),\,N(r_{f}R_{f}^{\circ},\,tB_{2}^{n})\Bigr\}\leqslant\exp\!\left(\tfrac{c_{1}\kappa_{n}^{2}n}{t^{2}}\right)$
for some $\kappa_{n}\geqslant 1$ and every $t\geqslant 1$, then
$N(g,t\odot f)\leqslant\exp\!\left(\frac{\kappa_{n}n}{t}\right)\qquad\text{for all }t\geqslant 1.$
@@
Let $t\geqslant 1$. If $x\in R_{f}$, then by definition $f(x)\geqslant\exp(-50n)f(0)$, and log-concavity gives
(3.19) $f(x/t)\geqslant f(x)^{1/t}f(0)^{1-1/t}=\left(\frac{f(x)}{f(0)}\right)^{1/t}f(0)\geqslant\exp(-50n/t)f(0)=\exp(-50n/t)$
because $f(0)=1$. Hence
@@
and Lemma 2.2 (iv) and (i) yield
(3.20) $N(g,t\odot f)\leqslant N(g,\exp(-50n/t)\,\mathds{1}_{R_{f}})=e^{50n/t}\,N(g,\mathds{1}_{R_{f}}).$
@@
(3.21) $g(x)=\sum_{k=0}^{\infty}g(x)\,\mathds{1}_{\{akr_{f}\leqslant|x|<a(k+1)r_{f}\}}(x)\leqslant\sum_{k=0}^{\infty}e^{-a^{2}k^{2}r_{f}^{2}/2}\,\mathds{1}_{a(k+1)r_{f}B_{2}^{n}}(x),$
@@
$\leqslant N\!\left(a(k+1)r_{f}B_{2}^{n},\,\tfrac{1}{\sqrt{\kappa_{n}t}}r_{f}B_{2}^{n}\right)\,N(r_{f}B_{2}^{n},\,\sqrt{\kappa_{n}t}R_{f}).$
@@
$N\!\left(a(k+1)r_{f}B_{2}^{n},\,\tfrac{1}{\sqrt{\kappa_{n}t}}r_{f}B_{2}^{n}\right)\leqslant\bigl(2\sqrt{\kappa_{n}t}\,a(k+1)+1\bigr)^{n},$
@@
$N(r_{f}B_{2}^{n},\sqrt{\kappa_{n}t}R_{f})\leqslant\exp\!\left(\frac{c_{1}\kappa_{n}n}{t}\right).$